• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name jpchen

    Log entry time 16:05:21 on March12,2007

    Entry number 195709

    keyword=Day Shift Summary

    Day shift summary:

    Shift workers: J. P. Chen(SL), C. Dutta(TO), S. Sirca.

    8:00 production on pb for kin06.
    8:20 scattering chamber vacuum rise to 2x10-5 and a spike to 5x10-5,then slowly back to below 1x10-5.
    9:05-9:11 right magnets frozen. stopped run 1446. rebooted ioc, NMR still frozen. rebooted ioc for NMR, all were fine.
    9:25 MCC called that they had ioc problem and will be no beam for ~15 minutes, stopped run 1447.
    9:57 beam back.
    run 1449 endrun became 1450 and no content.
    16:00 per Bob Michaels' request, stopped run 1458


    42 uA on Pb for kin06.
    Total charge from previous shift is 0.63C.
    run 1444, 7:39, 1M, 0.077C, total=0.707 C.
    run 1445, 8:13, 1M, 0.077C, total=0.784 C.
    run 1446, 8:46, 0.6M, 0.045C, total=0.829C.
    run 1447, 9:11, 0.4M, 0.032C, total=0.861C.
    run 1448, 9:57, 1M, 0.077C, total=0.938C.
    run 1449, 10:33, 1M, 0.077C, total=1.015C. No endrun.
    run 1450, 11:05, 1M, 0.077C, total=1.092C.
    run 1451, 11:46, 1M, 0.077C, total=1.169C.
    run 1452, 12:17, 1M, 0.077C, total=1.246C.
    run 1453, 12:51, 1M, 0.077C, total=1.323C.
    run 1454, 13:23, 1M, 0.077C, total=1.400C.
    run 1455, 13:57, 1M, 0.077C, total=1.477C.
    run 1456, 14:34, 1M, 0.077C, total=1.554C.
    run 1457, 15:06, 1M, 0.077C, total=1.631C.
    run 1458, 15:39, 0.7M, 0.05C, total=1.681C. No Endrun. Bob downloaded trigger after this run.